STAFF RESOURCES |
Additional Updates from the Billing Group Meeting. (Judy, Barry)
Policy Memo- discussion and clarification on: A new Replacement copy from the publisher, Binding/Damage Charge, No refunds on Long Overdue materials.
The group discussed RSG topic of requiring users to enter a Pin when placing Requests in I-Link. The ASC recommends requiring a pin.
Systems Upgrade-starting at 6pm Saturday 12/31/06-Sun 1/1/07
Testing-Known Issues for GL3.1 can be viewed at http://www.iris.rutgers.edu/docs/sirsidoc/relnotes/vgl31rn.pdf
J-Client training, our group will develop training for 3.1 in January with sessions held in February.
The Access Services & Systems Department Academic Calendar was reviewed and discussed
Discussion
Information Services, J Client
Circulation Group - Using Sakai to post announcements and email-announcements, although tests/quizzes are still problematic. They're testing GL3.1 and developing a standard toolbar for the access services service points.
Securities and Facilities Group - The November meeting has been cancelled. They're working with Anne Butman on setting up a PC Security Brown Bag Lunch.
Collection Management Group - The November meeting has been cancelled. They're working on procedures for IN-PROCESS/Notify items. They're planning for shifts at both the Chemistry and Physics Libraries.
Resource Sharing Group - The October meeting was held utilizing teleconferencing at 5 sites. They're working on testing GL3.1 and the Holdshelf Inventory reports are now running for most libraries.
Reserve Group - They're now using local Reserve-users for emailing for notification all reserve staff with one single email address. One message from a Faculty member/web page inquiry will get distributed to all Reserve staff at the unit. They tested a new scanner for Imaging Services. They have Sakai training session set up to train staff on how to link to Sakai links directly from E-reserve.
Billing Group-Working on Lost-Paid documentation
Student Supervisor's Group-no report
Judy - Meeting with RU-Student governments.
Michele - (via email) Interviewing for Dave Warner's position. They're sending over 100 boxes out to Better World Books. LSM has now weeded 34,127 items from the LSM collection. A shift in Bound Per is underway. New Stack Guidelines have been posted on the second floor.
Chris - 3 staff retired in TAS, Director's station training is progressing, the 180 day assumed lost notices for items due May 12, 2006 have gone out.
Penny - The Welcome desk is no longer staffed at the Douglass Library
Glenn - Posting on ILLIAD looking at Relay on Sirsi/Dynix
Katie - Shipping update on Camden to NB Campus mail
Jeff - It's been busy. They've been searching the items currently checked out to the CS users, in an attempt to clean up the records.
Roger - LSM UmdNJ line close to being filled. Krista is still covering at Physics.Bill and Kevin are still helping out at Kilmer. Ivy's Plus is scheduled for March 15/16 and RUL is hosting it.